### v1

# LAPD: Langevin-Assisted Bayesian Active Learning for Physical Discovery

Abstract
Discovering physical laws from data is a fundamental challenge in scientific research, particularly when high-quality data are scarce or costly to obtain. Traditional methods for identifying dynamical systems often struggle with noise sensitivity, inefficiency in data usage, and the inability to quantify uncertainty effectively. To address these challenges, we propose Langevin-Assisted Active Physical Discovery (LAPD), a Bayesian framework that integrates replica-exchange stochastic gradient Langevin Monte Carlo to simultaneously enable efficient system identification and robust uncertainty quantification (UQ). By balancing gradient-driven exploration in coefficient space and generating an ensemble of candidate models during exploitation, LAPD achieves reliable, uncertainty-aware identification with noisy data. In the face of data scarcity, the probabilistic foundation of LAPD further promotes the integration of active learning (AL) via a hybrid uncertainty-space-filling acquisition function. This strategy sequentially selects informative data to reduce data collection costs while maintaining accuracy. We evaluate LAPD on diverse nonlinear systems such as the Lotka-Volterra, Lorenz, Burgers, and Convection-Diffusion equations, demonstrating its robustness with noisy and limited data as well as superior uncertainty calibration compared to existing methods. The AL extension reduces the required measurements by around 60% for the Lotka-Volterra system and by around 40% for Burgers’ equation compared to random data sampling, highlighting its potential for resource-constrained experiments. Our framework establishes a scalable, uncertainty-aware methodology for data-efficient discovery of dynamical systems, with broad applicability to problems where high-fidelity data acquisition is prohibitively expensive.
1 Introduction
Data-driven discovery of dynamical systems has emerged as a transformative paradigm for extracting governing equations from observational data, enabling insights into complex phenomena in physics (Brunton et al.,, 2016), biology (Zheng et al.,, 2022) and engineering (Kaheman et al.,, 2020). Advances in sparse regression techniques, such as Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton et al.,, 2016; Rudy et al.,, 2016), and Neural Networks (Champion et al.,, 2019; Raissi et al.,, 2019), have demonstrated success in identifying parsimonious models under idealized conditions. Subsequently, several extensions have been developed to address real-world challenges (Messenger and Bortz,, 2021; Zhang and Lin,, 2018). However, real-world applications often face critical challenges: Experimental data are frequently corrupted by noise, scarce because of high acquisition costs, or sampled sparsely in high-dimensional domains. Traditional methods, which rely heavily on deterministic optimization or fixed regularization, struggle to quantify uncertainty in identified models, exhibit sensitivity to noise (Zheng and Lin,, 2024), and demand large datasets for reliable performance, which limits their utility in resource-constrained settings.
Bayesian methods, such as stochastic gradient Langevin dynamics (SGLD) and Hamiltonian Monte Carlo, leverage gradient-driven dynamics to explore posterior distributions of learned coefficients efficiently (Welling and Teh,, 2011; Neal et al.,, 2011). Although these methods improve traditional Markov Chain Monte Carlo (MCMC) techniques, they often face challenges when identifying convex dynamical systems with multi-modal regions, such as local mode trapping or sensitivity to identified parameters (Hirsh et al.,, 2022). To address this, replica exchange stochastic gradient Langevin dynamics (reSGLD) addresses these limitations by combining parallel sampling chains at multiple temperatures (Deng et al.,, 2020; Zheng et al.,, 2024), which enables aggressive exploration of identified coefficient space through high-temperature chains while low-temperature chains refine high-probability regions. This balance is critical for identifying sparse, interpretable dynamical systems from noisy data and providing robust UQ.
In resource-constrained scenarios, AL strategies minimize data acquisition costs by iteratively selecting measurements that maximize information gain. Traditional approaches prioritize either uncertainty reduction (e.g., querying regions with high predictive variance) or space-filling designs (Meng and Karniadakis,, 2020) (e.g., uniform domain coverage). However, purely uncertainty-driven physical discovery risks over-exploiting localized regions, while diversity-focused strategies neglect model confidence. Hybrid acquisition functions, which combine uncertainty and spatial diversity metrics, mitigate this trade-off by selecting points that resolve model ambiguity while ensuring broad coverage of the system’s phase space. For dynamical systems, this might involve targeting states where trajectory predictions diverge across posterior samples or where sparse measurements leave governing equations underdetermined. Bayesian frameworks naturally enable such strategies by quantifying epistemic uncertainty, yet existing methods often lack scalable sampling or adaptive acquisition protocols.
To address such issues, the proposed Langevin-Assisted Active Physical Discovery (LAPD) unifies these advances into a cohesive framework for data-efficient, uncertainty-aware system identification. By balancing exploration and exploitation with reSGLD and a hybrid uncertainty-space-filling acquisition function in AL, LAPD achieves robust posterior estimation even with noisy and limited data. The framework adaptively prioritizes measurements that maximally reduce uncertainty in model coefficients while ensuring diverse sampling of the dynamical landscape—effectively minimizing experimental costs. Our contributions include:
  • a scalable Bayesian framework considering reSGLD for uncertainty-aware system identification;
  • an AL protocol that hybridizes uncertainty and diversity criteria for efficient data sampling;
  • empirical validation across the Lotka-Volterra system, Lorenz system, Burgers’ equation, and convection-diffusion system to validate its accurate UQ and data acquisition.
This work bridges the gap between probabilistic inference and data efficiency, which offers a principled tool for scenarios where high-fidelity data are prohibitively expensive.
The remainder of this paper is organized as follows. Section 2 formulates the problem of data-driven dynamical system discovery. Section 3 introduces the LAPD framework, which details its integration of reSGLD for uncertainty-aware model identification. Section 4 validates the framework through experiments on four non-linear systems: the Lotka-Volterra system, the Lorenz system, the Burgers equation, and the convection-diffusion equation, which demonstrates robustness to noise and data scarcity. In Section 5, we extend LAPD with an AL strategy that leverages its UQ capabilities to optimize data acquisition, with benchmarks on the Lotka-Volterra and Burgers’ systems under limited-data regimes. Finally, Section 6 discusses the broader implications, limitations, and future directions.
2 Problem Statement
2.1 Physical Discovery
We study a class of differential equations given by:
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbf{u}(t)=\mathbf{f}(\mathbf{u}(t),t),$ (1)
where the state vector is defined as $\mathbf{u}(t)=[u_{1}(t),u_{2}(t),\cdots,u_{d}(t)]^{\intercal}\in\mathbb{R}^{d}$, and $t$ denotes the corresponding temporal inputs. For simplicity, we exclude spatial inputs from the formulation (1), although such extensions could include partial derivatives with respect to spatial variables. The dynamics of the system is encoded by the function $\mathbf{f}(\cdot)$, which is unknown and usually of non-linear form.
The dynamics of the system $\mathbf{f}(\cdot)$ is learned from the data collected at time $t=t_{n},\ n=1,2,3,\cdots,N$. The size of the data collected is $N$, with the collected data denoted as $\mathbf{U}=[\mathbf{u}(t_{1}),\mathbf{u}(t_{2}),\cdots,\mathbf{u}(t_{N})]^{\intercal}\in\mathbb{R}^{N\times d}$. The learning of $\mathbf{f}(\cdot)$ is accomplished by assuming that the underlying dynamical system can be expressed as a linear combination of functions from a candidate library $\bm{\Theta}(\mathbf{U})=\begin{bmatrix}\theta_{1}(\mathbf{U}),\cdots,\theta_{m}(\mathbf{U})\end{bmatrix}\in\mathbb{R}^{N\times m}$:
$\dot{\mathbf{U}}=\bm{\Theta}(\mathbf{U})\bm{\Xi},$ (2)
where $\dot{\mathbf{U}}\in\mathbb{R}^{N\times d}$ denotes the temporal derivative of the $\mathbf{U}$, $\bm{\Xi}\in\mathbb{R}^{m\times d}$ are coefficients that determine which specific functions from the library $\bm{\Theta}(\mathbf{U})$ contribute to the system dynamics. A typical choice for the candidate functions includes polynomials of the state variables (as well as the partial derivatives of $x$ for PDEs), as they are commonly associated with canonical models of dynamical systems. Fourier-based libraries, incorporating terms such as $sin(\mathbf{x})$ and $cos(\mathbf{x})$, are also frequently employed.
A key assumption to characterize the system is that the coefficient matrix $\mathbf{\Xi}$ is sparse, i.e., the system can be represented by a small number of candidates in the library $\bm{\Theta}$. The standard Sparse Identification of Nonlinear Dynamics (SINDy) (Brunton et al.,, 2016; Rudy et al.,, 2016) uses a sequentially thresholded least-squares (STLSQ) algorithm to perform sparse inference. The STLSQ method serves as a proxy for $l_{0}$ optimization and comes with established convergence guarantees (Zhang and Schaeffer,, 2018). This algorithm combines sparse regression techniques like Ridge and Lasso regression with thresholding to eliminate insignificant terms. To enhance robustness and incorporate UQ, we reformulate the identification of system (1) within a Bayesian framework as follows.
2.2 Bayesian Inference for Data-driven Discovery
A Bayesian framework can be applied to identify and estimate the uncertainty of the systems (1), or specifically, we aim to infer the posterior distribution of $\mathbf{\Xi}$. Let $p(\mathbf{\Xi})$ denote the prior distribution and $p(\mathbf{U}|\mathbf{\Xi})$ the likelihood of the observed data $\mathbf{U}$ given the system parameterized by $\mathbf{\Xi}$. Following the Bayes formula, the posterior distribution of $\mathbf{\Xi}$ given the $N$ collected data snapshots $\mathbf{U}$ is:
$p(\mathbf{\Xi,\Sigma}|\mathbf{U})\propto p(\mathbf{\Xi})p(\mathbf{\Sigma})\prod_{i=1}^{N}p(\mathbf{u}_{i}|\mathbf{\Xi,\Sigma}),$ (3)
where the term $\mathbf{\Sigma}$ encompasses additional parameters of the probabilistic model, including variables such as the standard deviation associated with the random noise term. The computation of the posterior distribution (Equation (3)) typically presents analytical intractability. In such cases, sampling-based approaches, such as Markov Chain Monte Carlo (MCMC) methods, provide a viable solution. The approximated posterior distribution enables model reconstructions and trajectory forecasts given observational data, allowing quantitative uncertainty assessment of model coefficients and predicted trajectories.
Within the Bayesian framework, it necessitates a prior distribution $p(\mathbf{\Xi})$ that promotes sparsity while robustly handling coefficients of varying magnitudes. The Regularized Horseshoe prior (Piironen and Vehtari,, 2017) is particularly well-suited for this task, as it combines strong shrinkage for irrelevant coefficients with controlled regularization for significant ones—avoiding over-shrinkage of true signals common in standard sparsity priors. To adopt this, we first assign each coefficient $\beta_{i}$ a local shrinkage parameter $\tilde{\lambda}_{i}$, which determines how strongly it is shrunk toward zero:
$\beta_{i}|\tilde{\lambda}_{i},\tau,c\sim\mathcal{N}(0,\tilde{\lambda}_{i}^{2}\tau^{2}),\ \ \text{where}\ \ \tilde{\lambda}_{i}=\frac{c\lambda_{i}}{\sqrt{c^{2}+\tau^{2}\lambda_{i}^{2}}}.$
[…]
With the Regularized Horseshoe prior, most identified coefficients are shrunk aggressively toward zero, which effectively removes irrelevant or redundant basis functions. On the other hand, significant coefficients escape shrinkage due to the slab, which preserves their magnitude. The inverse-Gamma prior on $c^{2}$ ensures that the slab variance remains bounded, which avoids failed coefficient identification in MCMC sampling. We will elaborate on how we employ gradient-based MCMC methods in both identifying system coefficients and provide robust UQ within limited data in the next section.
3 Langevin-Assisted Active Physical Discovery
In this work, we introduce LAPD, which incorporates Langevin MCMC methods into the discovery of physical laws in a Bayesian way via a sequential threshold method. LAPD utilizes the stochastic gradient in the situation of sufficient data collection to improve efficiency and can be exploited for AL in the face of data scarcity (i.e., the time derivative calculation or collection is expensive). We start with the data collected $\mathbf{U}\in\mathcal{R}^{N\times d}$ from the ODE or PDE systems as described before, with the task of identifying the system in terms of (1) and determining the posterior distribution of the model coefficient $\mathbf{\Xi}\in\mathbb{R}^{k\times d}$, where $k$ is the number of candidates in each row of the library $\mathbf{\Theta}(\mathbf{U})$.
3.1 Uncertainty Quantification with Langevin MCMC
Langevin-type algorithms are grounded in Langevin diffusion, a stochastic process governed by the stochastic differential equation (SDE):
$d\mathbf{\Xi}_{t}=-\nabla U(\mathbf{\Xi}_{t})\,dt+\sqrt{2\tau}\,d\mathbf{W}_{t},$ (4)
where $U(\cdot)$ represents the energy function, $\{\mathbf{W}_{t}\mid t\geq 0\}$ denotes the standard Brownian motion on $\mathbb{R}^{m\times d}$, and $\tau>0$ is the temperature parameter. Under mild conditions on $U$, it is well-established that the diffusion process described by (4) admits a unique strong solution $\{\mathbf{\Xi}_{t},t\geq 0\}$, which is a Markov process. Moreover, as $t\to\infty$, the distribution of $\mathbf{\Xi}_{t}$ converges to the invariant distribution $\pi_{\tau}$, characterized by the density $\pi_{\tau}(\mathbf{\Xi})\propto\exp(-U(\mathbf{\Xi})/\tau)$.
To numerically approximate Langevin diffusion (4), the forward Euler discretization is commonly employed. This leads to the iterative update:
$\tilde{\mathbf{\Xi}}_{k+1}=\tilde{\mathbf{\Xi}}_{k}-\eta_{k}\nabla U(\tilde{\mathbf{\Xi}}_{k})+\sqrt{2\eta_{k}\tau}\,\zeta_{k},$ (5)
where $\zeta_{k}\sim\mathcal{N}(0,I_{md\times md})$ represents independent Gaussian noise at each iteration, and $\eta_{k}$ is the step size at iteration $k$. To enable scalable and efficient sampling while maintaining convergence to the target distribution under appropriate conditions, we further consider its stochastic gradient version for efficient large-scale data applications. Specifically, the energy function is approximated with a mini-batch of data $\mathbf{B}$ uniformly subsampled from the given data (Welling and Teh,, 2011):
$\tilde{L}(\tilde{\mathbf{\Xi}})=-\log p(\tilde{\mathbf{\Xi}})-\frac{N}{|\mathbf{B}|}\sum_{\mathbf{x}_{i}\in\mathbf{B}}\log P(\mathbf{x}_{i}\mid\tilde{\mathbf{\Xi}}),$ (6)
Given the problem setup (2), the gradient $\nabla\tilde{L}(\tilde{\mathbf{\Xi}})$ is calculated as
$\nabla\tilde{L}(\tilde{\mathbf{\Xi}})=-\nabla\log p(\tilde{\mathbf{\Xi}})-\frac{N}{|\mathbf{B}|}(\mathbf{\Theta}(\mathbf{B})^{\intercal}(\dot{\mathbf{B}}-\mathbf{\Theta}(\mathbf{B})\tilde{\mathbf{\Xi}})).$ (7)
For an appropriately chosen step size schedule $\eta_{k}$, it has been established that the Langevin MCMC converges to the target stationary distribution (Durmus and Moulines,, 2018). Despite the theoretical guarantees, the existing non-asymptotic convergence results can be difficult to interpret and apply in practice, particularly when step size schedules must balance convergence speed with accuracy. To address these challenges, several methods, such as underdamped Langevin MCMC (Cheng et al.,, 2018), Hamiltonian Monte Carlo (Neal et al.,, 2011), cyclical SGMCMC (Zhang et al.,, 2019), reSGLD (Chen et al.,, 2018), etc., have been developed as extensions of the Langevin MCMC framework. These methods aim to enhance computational efficiency and scalability, especially in high-dimensional settings.
We consider reSGLD to balance exploration and exploitation when identifying system coefficients, which offers a blend of efficiency and robustness in traversing different energy levels to address large-scale non-convex sampling problems (Deng et al.,, 2020; Zheng et al.,, 2024). It simulates a high-temperature chain for exploration and a low-temperature chain for exploitation. The sampling process is shown as follows:
$\tilde{\mathbf{\Xi}}_{k+1}^{(1)}$ $=\tilde{\mathbf{\Xi}}_{k}^{(1)}-\eta_{k}\nabla\tilde{L}(\tilde{\mathbf{\Xi}}_{k}^{(1)})+\sqrt{2\eta_{k}\tau_{1}}{\zeta}_{k}^{(1)}$ (8)
$\tilde{\mathbf{\Xi}}_{k+1}^{(2)}$ $=\tilde{\mathbf{\Xi}}_{k}^{(2)}-\eta_{k}\nabla\tilde{L}(\tilde{\mathbf{\Xi}}_{k}^{(2)})+\sqrt{2\eta_{k}\tau_{2}}{\zeta}_{k}^{(2)},$
where $\tilde{\mathbf{\Xi}}_{k}^{(1)}$ and $\tilde{\mathbf{\Xi}}_{k}^{(2)}$ denote the sampling results of two chains with temperatures $\tau_{2}>\tau_{1}$. Furthermore, we swap the Markov chains in (8) with the corrected swapping rate for mini-batch setting:
$\tilde{S}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(1)},\tilde{\mathbf{\Xi}}_{k}^{(2)}\big{)}=e^{\left(1/{\tau_{1}}-1/{\tau_{2}}\right)\left(\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(1)}\big{)}-\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(2)}\big{)}-\left(1/{\tau_{1}}-1/{\tau_{2}}\right)\frac{\tilde{\sigma}^{2}}{C}\right)}$ (9)
where $\tilde{\sigma}^{2}$ approximates the variance of $\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(1)}\big{)}-\tilde{L}\big{(}\tilde{\mathbf{\Xi}}_{k}^{(2)}\big{)}$ and $C$ acts as an adjustment to balance acceleration and bias. The algorithm proceeds iteratively until a specified number of iterations is reached. Generate uniform $u\in[0,1]$, if $u<\tilde{S}$, then perform the swapping between two chains. The parameter sets $\{\tilde{\mathbf{\Xi}}_{k}^{(1)}\}_{k=1}^{K+1}$ are produced as outputs for analytical purposes.
3.2 Bayesian Physical Discovery with Limited Data
The integration of AL into data driven discovery of dynamical systems attempts to address two fundamental challenges inherent to learning parsimonious dynamical models from limited and costly data. First, accurate estimation of state derivatives $\mathbf{\dot{U}}$, which is critical for regression-based identification of (1), becomes unreliable when the data are irregularly sampled or sparsely distributed. Traditional numerical differentiation (e.g., finite differences) amplifies noise and errors under these conditions, particularly when time intervals are large or nonuniform. Second, dynamical systems often exhibit heterogeneous informativity across their state space: sparse governing terms dominate in specific regimes (e.g., near bifurcations or transient states), while other regions contribute minimally to identifying the dynamical systems. Passive or uniform sampling risks undersampling these critical regions, leading to biased or even incorrect models.
AL directly targets these issues by prioritizing data acquisition, where the uncertainty in the learned dynamics is maximized. By tackling (2) within a Bayesian framework, posterior uncertainties over the coefficients $\mathbf{\Xi}$ in (2) can be quantified. Regions of high predictive entropy or variance in $\mathbf{\dot{U}}$ correspond to states in which the model cannot confidently distinguish between candidate terms in the library $\mathbf{\Theta}(\mathbf{U})$. Actively querying derivatives in these states maximizes the information gain on $\mathbf{\Xi}$, which efficiently resolves ambiguities in the sparsity pattern while minimizing redundant sampling (Riis et al.,, 2023; Pickering et al.,, 2022; Gramacy and Lee,, 2009; Fasel et al.,, 2022).
To formulate AL, we first denote $\mathcal{D}=\{\mathbf{U},\dot{\mathbf{U}}\}$. With the information of the posterior $p(\bm{\Xi}|\mathcal{D})$,
[…]
During implementation, we estimate $\mathbb{E}_{p(\boldsymbol{\bm{\Xi}}\mid\mathcal{D})}\left[\sigma_{\boldsymbol{\bm{\Xi}}}^{2}(\dot{\mathbf{u}})\mid\boldsymbol{\bm{\Xi}}\right]$ with the predictive variance $\sigma^{2}(\dot{\mathbf{u}})$. The higher the variance, the more uncertainty we have for the sample.
We sample $P$ groups of the posterior distribution of model parameters and calculate $\dot{\mathbf{u}}^{pred}_{i}$ for each model $i$. The predictive variance is calculated as
$\sigma^{2}({\dot{\mathbf{u}}})=\frac{1}{P}\sum_{i=1}^{P}(\dot{\mathbf{u}}^{pred}_{i}-\bar{\dot{\mathbf{u}}}^{pred})^{2},$
leading to the selection of potential design points with high uncertainty. Through experiments, it is observed that using the uncertainty acquisition function alone may cause the selected points to cluster in different rounds. While those points are informative, we are missing information from other regions, which causes a rather biased model favoring and overfitting one region of the data.
To solve this problem, we combine the above-mentioned uncertainty acquisition function with a space-filling design to spread the design points. Here we adopt the maximin distance criterion (Loeppky et al.,, 2010; Yu and Kim,, 2010), which selects the next design points that maximize the minimum distance to the current design points. Here, the maximin criterion is adjusted by the density of the data.
$d(\mathbf{\mathbf{\Theta}(\mathbf{u})})=\text{min}||\mathbf{\Theta}(\mathbf{u})-\mathbf{\Theta}(\mathbf{u}_{i})||_{2}\cdot\text{density}(\mathbf{u})^{\lambda},\text{for all}~{}\mathbf{u}_{i}\in\mathcal{M},$ (10)
where $\lambda\geq 0$ is a tunable hyperparameter that controls how much density penalizes on the distance. when $\lambda=0$, the criterion returns to the maximin criterion with no adjustment by density. Density-based adjustment has been adopted in AL research for both classification and regression tasks to achieve different goals, such as increasing the representativeness of the data (Wang et al.,, 2021), mitigating the sensitivity of the Euclidean distance to the scale and distribution of the data (Donmez and Carbonell,, 2007) and preventing outliers added to the training data (Zhu et al.,, 2008). The density measure includes KDE, cosine distance-based measurement, and K-Nearest-Neighbor-based density measure. We opt for the K-Nearest-Neighbor-based density measure over KDE or cosine similarity because KNN offers computational simplicity and directly captures local data density without making strong assumptions about the underlying data distribution, and unlike cosine similarity, it preserves both directional and magnitude information that is essential for identifying truly representative samples in our feature space. The K-Nearest-Neighbor-based density is calculated as the inverse of one point’s average distance to its K nearest neighbor.Our goal here is to penalize on data whose density is low, since in this situation, there is a higher chance that it holds high maximin distance with data already selected, especially when the data itself is also of high magnitude. In the sense that without density adjustment, data of low density and high magnitude are preferred, while data with high density and high magnitude are disregarded. This prevents the criterion from achieving the goal of space-filling and potentially favoring the same group of data as the uncertainty criterion. When the data distribution is quite uniform or the magnitude does not vary too much, there is no need to do a density adjustment.
Two criterion are standardized into $[0,1]$ to get $\tilde{\sigma^{2}}({\dot{\mathbf{u}}})$ and $\tilde{d}(\mathbf{\mathbf{\Theta}(\mathbf{u})})$ and the new acquisition function is defined by weighted sum, with $\alpha$ being the tunable weight,
$C(\mathbf{u})=\alpha\,\tilde{\sigma^{2}}({\dot{\mathbf{u}}})+(1-\alpha)\,\tilde{d}(\mathbf{\mathbf{\Theta}(\mathbf{u})}).$ (11)
Input: Pool of potential design points $\mathcal{D}$; number of initial points $n$; batch size of selected points per round $m$, maximum allowed sample size $N_{max}$, tolerance of convergernce Tol.
Output: Stabilized model with coefficients $\mathbf{\Xi}$
Randomly select $n$ points, $\mathbf{u}_{1},\dots,\mathbf{u}_{n}\in\mathbf{U}$ with derivatives $\dot{\mathbf{u}}_{1},\dots,\dot{\mathbf{u}}_{n}$ available;
Set of selected points =$\mathcal{M}=\{(\mathbf{u}_{1},\dot{\mathbf{u}}_{1}),\dots,(\mathbf{u}_{n},\dot{\mathbf{u}}_{n})\}$, $\mathcal{D}\leftarrow\mathcal{D}/\mathcal{M}$;
Initialize $\mathcal{E}_{0}$ as a very large number;
while $|\mathcal{M}|\leq N_{max}$ do
       Use $\mathcal{M}$ points to construct the library $\mathbf{\Theta}$ and perform LAPD to estimate the posterior distribution $p(\boldsymbol{\bm{\Xi}}\mid\mathcal{M})$;
       for all potential design points $\mathbf{u}\in\mathcal{D}$ do
             Calculate $C(\mathbf{u})$
       end for
      for $i=1,\dots,m$ do
            Select the sample $\mathbf{u}$ with highest $C(\mathbf{u})$;
             Collect additional data and compute derivative $\mathbf{\dot{u}}$ if not available;
             $\mathcal{M}\leftarrow\mathcal{M}\cup{(\mathbf{u},\dot{\mathbf{u}})},\mathcal{D}\leftarrow\mathcal{D}/\mathcal{M}$
       end for
      Calculate $\mathcal{E}$;
       if $\frac{|\mathcal{E}-\mathcal{E}_{0}|}{\mathcal{E}}<\text{Tol}$ then
            break;
       end if
      $\mathcal{E}_{0}\leftarrow\mathcal{E}$;
end while
Algorithm 1 Active Learning in LAPD: Iteratively improve the model by selecting design points that maximize the acquisition function, updating the posterior, and retraining the model until convergence.
Intuitively, the synergy arises from AL’s ability to exploit the structured sparsity of dynamical systems. Unlike generic regression tasks, SINDy can identify a small subset of relevant terms from a potentially large library. Correlated or redundant terms create degeneracies that passive data may fail to disentangle. AL breaks these degeneracies by selecting states in which the predictions of competing sparse models diverge most significantly. To make it clearer, in a system with cubic nonlinearity terms, this strategy will favor sampling in high-energy regions where the cubic terms dominate rather than in low-energy regions where the linear-term approximations suffice. It not only reduces the number of required $\mathbf{\dot{U}}$ evaluations but also enhances robustness to noise and model misspecification, which ensures that the identified dynamics are both parsimonious and physically consistent.
3.3 The Proposed Algorithms
We now introduce the procedure for implementing LAPD. This framework can be applied to scenarios with sufficient data, whether clean or noisy, utilizing reSGLD for parameter sampling, and it can also be applied to scenarios with limited data where AL will help to minimize data acquisition cost. The proposed LAPD framework is summarized in Figure 1.
Figure 1: The LAPD framework is demonstrated through the identification of the Lotka-Volterra system. The process begins with Step 1: Data Preparation, where the dataset $\mathbf{U}$ is obtained from the differential equations under investigation. Step 2: Sampling: The (stochastic) gradient $\tilde{L}(\tilde{\mathbf{\Xi}}_{k})$ is computed as outlined in the Gradient Calculation procedure (bottom). Step 3: Sequential Thresholding: Based on the sampling outcome, thresholding is applied by removing bases with small absolute coefficients from the library; for example, as in the plot, the $xy$ basis for state $x$ is excluded by thresholding. If the retained library differs from the previous iteration, another sampling round is conducted. Step 4: Active Learning: If AL is adopted, continue to add new training points according to the hybrid acquisition function and restart the training. Step 5 Reconstruction: Once the final posterior distribution $P(\tilde{\mathbf{\Xi}}|\mathbf{U})$ is obtained, the system is reconstructed with uncertainty quantification.
Data preparation. To begin, prepare the time derivatives $\dot{\mathbf{U}}$ and the library $\bm{\Theta}(\mathbf{U})$. If $\dot{\mathbf{U}}$ is not directly observed, its calculation depends on the type of data. The Finite Difference in PySINDy is adopted as the derivative calculation method for clean or noisy data. For ODEs, the library is made up of polynomials of the system states. For PDEs, partial derivatives of the position $x$ will need to be calculated before constructing the library, where the finite difference is also adopted. We also attempted other approaches to approximate the derivative terms in the face of noise, like spline or Savitzky-Galoy, and found that Finite Difference best suits the proposed framework considering approximation efficiency, accuracy, and robustness.
Parameter Sampling. The posterior sampling process involves several key steps. First, select the appropriate priors for the coefficients. Next, define the posterior distribution using the negative log-posterior (6) as the energy function. To ensure that the stationary distribution of the sampler is proportional to $\exp(-L(\mathbf{\Xi}))$, set the posterior as untempered with $\tau=1$. Additionally, apply a Metropolis-Hastings (MH) adjustment to correct any bias introduced during sampling. After defining these parameters, set an appropriate burn-in period. Discard the samples from the burn-in phase and use the remaining samples to estimate the posterior distribution of the model parameters. Sequential thresholding is employed after posterior samples of the coefficients following (8), where the coefficients with posterior modes lower than some threshold $c$ are removed from the basis (Brunton et al.,, 2016), which helps to achieve sparse results.
Data Acquisition for Active Learning. In the face of data scarcity, if AL is adopted, new potential design points will be added to the selected points $\mathcal{M}$. The implementation of AL in LAPD is shown in Algorithm 1. Return to Step 1 for new data preparation, for example, update the derivative and the basis library with new $\mathcal{M}$. If model training is completed, go on to reconstruct the model and estimate the uncertainty.
Model Reconstruction and Uncertainty Quantification. For ordinary differential equation (ODE) systems, predictions are made by solving the initial value problem. The solution involves integrating over the posterior distribution of $\tilde{\mathbf{\Xi}}$ using the formula:
[…]
In real applications, it is often not feasible to solve the integral in the prediction step analytically. Instead, numerical integration can be used to compute the system’s trajectory, leveraging samples of $\tilde{\mathbf{\Xi}}$ obtained from the posterior distribution. For partial differential equation (PDE) systems, given $\tilde{\mathbf{\Xi}}$, the prediction is either made by solving the system numerically, i.e., converting the system to ODE through Fast Fourier Transform (FFT), or by solving it analytically.
UQ for the coefficients comes from the samples, where the estimated distribution can be learned. UQ for the trajectory is achieved by bootstrapping from the posterior coefficient samples and make prediction with each sample.
4 Numerical Experiments with Langevin-Assisted Physical Discovery
In this part, we conduct a total of four experiments: two for ODE systems (Lotka-Volterra and Lorenz) and two for PDE systems (Burgers’ Equation and Convection-Diffusion Equation). The baseline models used are SINDy (or PDE-FIND for PDEs) and UQ-SINDy. SINDy, a frequentist method, is less robust to noise compared to LAPD, as we will demonstrate later. UQ-SINDy, a Bayesian approach, provides UQ and identifies sparse systems. However, its reliance on sparsity-inducing priors can lead to failure in discovering the true system when the prior is not perfectly chosen. Additionally, we compare different Langevin MCMC methods within LAPD, including SGLD, cyclical SGMCMC, and reSGLD.
Cyclical SGMCMC follows the same iteration formula as (5), but the step size is adjusted every round to balance between exploration and sampling. The stepsize at iteration $k$ is defined as
$\eta_{k}=\frac{\eta_{0}}{2}\left[\cos\left(\pi\frac{\text{mod}(k-1,\lfloor K/M\rfloor)}{\lfloor K/M\rfloor}\right)+1\right],$
where $\eta_{0}$ is the initial stepsize, $M$ is the number of cycles, and $K$ is the number of total iterations. During implementation of Langevin MCMC methods, the Metropolis-Hastings step is carried out to ensure that the distribution of the iterated instances $\mathbf{\Xi}_{k}$ sampled at round $k$ converges to the correct distribution $\Pi$ when $k\rightarrow\infty$ (Dwivedi et al.,, 2019; Chewi et al.,, 2021). Unless otherwise specified, the default sampling method adopted in this work is the 2-chain reSGLD, and the prior adopted is the regulized horseshoe with $\nu=4,s=2$ as in (Hirsh et al.,, 2022).
Before conducting the experiments, we outlined the setup details. The noise-free data of ODEs is simulated numerically with solve_ivp method from the Scipy package in Python using the ’RK45’ integration method, setting 1e-3 for relative toerance and 1e-6 for absolute tolerance. The 1d Burgers’ system was converted into a system of ODEs using FFT, and the odeint function from scipy.integrate was used to numerically generate noise free data this system. The Convection-Diffusion system was solved analytically in python with Numpy. The noisy data is generated by adding independent and identically distributed (i.i.d.) Gaussian white noise with zero mean to the clean data. The variance of the noise for each dimension of the state is calculated as $\sigma=p\cdot\text{std}(\mathbf{x})$, where $\text{std}(\mathbf{x})$ is the standard deviation of the clean data along that dimension, and $p$ is the chosen noise level (e.g., 5%). All models are trained with the same derivate $\dot{\mathbf{U}}$ and library $\mathbf{\Theta}(\mathbf{U})$ prepared, and the same threshold if sequential thresholding adopted, making sure fair comparison.
We used three error metrics to evaluate the performance of different models, namely, Error Bar, Mean Squared Error (MSE) and Akaike Information Criterion (AIC). Error Bar is uncertainty-aware and works well in the Bayesian setting. The Error Bar criterion is defined as below following Zhang and Lin, (2018):
$\mathcal{E}(\tilde{\mathbf{\Xi}})=\sum_{\begin{subarray}{c}i=1\\ \tilde{\Xi}_{i}\neq 0\end{subarray}}^{md}\frac{\tilde{s}_{i}^{2}}{\tilde{\Xi}_{i}^{2}},$ (12)
with $md$ being the total number of entries in the parameter, $\tilde{s}_{i}$ representing the estimated variance, and $\tilde{\Xi}_{i}$ denoting the estimated posterior mode. The posterior samples generated from the MCMC method allow direct calculation of the estimated $\tilde{s}_{i}$ and $\tilde{\Xi}_{i}$. If $\tilde{\Xi}_{i}=0$, this means that at location $i$, the estimated mode is equal to zero, or the estimated mode is within the threshold and therefore set to zero, then the corresponding basis would be eliminated from the library for the corresponding state. The Bayesian error bar works both in model comparison and in threshold selection. The example of threshold selection is given in Section 4.4. A lower Error Bar indicates a more reliable coefficient result. MSE quantifies the discrepancy between the predicted trajectory of the identified model and the actual measurements, which is calculated as
$MSE=\frac{1}{Nd}\sum_{i=1}^{N}\sum_{j=1}^{d}(\mathbf{U}_{ij}-\tilde{\mathbf{U}}_{ij})^{2},$
where $\mathbf{U}$ denotes the true trejectory data (noise-free) and $\tilde{\mathbf{U}}$ denotes the predicted trajectory data, $N$ and $d$ representing the number of collection and state dimension, respectively. The smaller the MSE, the better the model learned fits the data. AIC deals with the trade-off between the goodness of fit of the model and the simplicity of the model. Here, AIC can be calculated based on the MSE,
$AIC=2k+Nd\,\ln(\text{MSE}),$
where $k$ denotes the number of non-zero coefficients. The smaller the AIC, the more favorable the model, as it effectively balances model fit and parsimony, penalizing unnecessary complexity while rewarding accuracy.
4.1 Lotka-Volterra system
We begin by assessing the method using the Lotka-Volterra system, often known as the predator-prey model. This system is extensively utilized to depict the interactions between two competing populations. Initially formulated by Lotka to simulate chemical reactions, it has since become a foundational framework for analyzing dynamics in biological systems and economic models. The Lotka-Volterra system can be characterized by the following nonlinear ODE system:
$\begin{cases}\dot{x}=\alpha x-\beta xy,\\ \dot{y}=-\delta y+\gamma xy.\end{cases}$ (13)
In this experiment, we simulate the data with the initial condition $[x_{0},y_{0}]=[10,5]$, and the system parameters $\alpha=1.0,\beta=0.1,\delta=1.5,\gamma=0.075$. Starting from $t=0$, $5,000$ time steps of states are simulated, with $\Delta t=5\cdot 10^{-3}$. The library consists of polynomial terms up to order 2, with the library basis being $[1,x(t),y(t),x(t)^{2},x(t)y(t),y(t)^{2}]$. The collected data are then corrupted with additive noise 5%.
The coefficients learned from various models are presented in Table LABEL:lotka_table. The results indicate that LAPD performs well in accurately identifying the true system structure, with no extraneous or missing terms, whereas SINDy and UQ-SINDy introduce additional terms not present in the true system. LAPD also shows better performance compared to other baseline methods, as reflected in lower MSE and AIC values. UQ-SINDy, which also uses a Bayesian approach to estimate the posterior distribution of the coefficients, relies on sparsity-inducing priors such as the horseshoe prior to exclude extraneous terms. However, this method is sensitive to the choice of prior and often does not achieve the desired system sparsity.
Figure 2(a) illustrates the in-sample predictions using coefficients learned from 10 different groups of priors with varying coefficients for the horseshoe prior. With $\tau_{0}\sim U(0.05,0.3)$, $\nu\sim U(0,5)$, $s\sim U(0,3)$. The plot shows considerable variability in the predictions of UQ-SINDy, with some groups failing to capture the state fluctuations. In contrast, the prediction trajectories generated by LAPD closely follow the true trajectory. This robustness is mainly due to the fact that LAPD’s coefficient updates are primarily driven by likelihood rather than the specific choice of prior, reducing sensitivity to prior selection. Additionally, the use of sequential thresholding helps achieve system sparsity rather than relying merely on sparsity-inducing priors.
Figure 2(b) provides 95% confidence intervals in the prediction space. The uncertainty coverage effectively captures the true trajectory of the Lotka-Volterra system, particularly in cases where the model’s predictions deviate from the truth. When the model predictions align well with the true trajectory, the uncertainty coverage remains appropriately modest. This highlights the reliability of the LAPD in both accurate prediction and uncertainty estimation.
Figure 2: (a) Predicted trajectories of UQ-SINDy and LAPD on the Lotka-Volterra system. Gray dots represent training data, purple dashed lines are trajectories generated by UQ-SINDy, and green solid lines are yielded by LAPD. Both UQ-SINDy and LAPD are run 10 times with different prior initializations. (b) Trajectory prediction and confidence interval with LAPD on Lotka-Volterra system, with the 95% confidence interval. The dashed vertical line marks the transition from in-sample prediction to out-of-sample prediction.
Table 1: Learned coefficients with different models on Lotka Volterra System with 5% noise
Basis True SINDy UQ-SINDy LAPD
X Y X Y X Y X Y
1 - - 4.158e-1 - 1.814e-2 -4.512e-2 - -
x 1.000e0 - 9.858e-1 - 9.613e-1 -2.194e-2 9.966e-1 -
y - -1.500e0 - -1.428e0 1.793e-2 -1.301e0 - -1.469e0
$x^{2}$ 3.281e-4 9.838e-4 - -
$xy$1.000e-1 7.500e-2 -1.000e-1 7.446e-2 -9.707e-2 6.935e-2 -1.000e-1 7.386e-2
$y^{2}$1.957e-3 -4.288e-3 - -
Error Bar - - 1.017e3 1.747e-3
MSE - 4.621 19.541 2.926
AIC - 7.663e3 1.508e4 5.377e3
4.2 Lorenz System
Consider Lorenz system of the form:
$\begin{cases}\dot{x}=\sigma(y-x),\\ \dot{y}=x(\rho-z)-y,\\ \dot{z}=xy-\beta z,\end{cases}$ (14)
with the system dynamics being governed by three parameters: the Prandtl number $\sigma$, the Rayleigh number $\rho$, and the aspect ratio $\beta$. This system is renowned for its chaotic behavior and is widely used to study complex, nonlinear dynamics in fields such as meteorology, fluid dynamics, and climate modeling. The interplay between these parameters gives rise to intricate and often unpredictable trajectories, making the Lorenz system a classic example of deterministic chaos. We simulate the system with initial conditions $\begin{bmatrix}x(0),y(0),z(0)\end{bmatrix}=\begin{bmatrix}-8,8,27\end{bmatrix}$, starting from $t=0$ and end at $t=5$ with time intervals $\Delta t=1\cdot 10^{-3}$. Here we consider the library consisted of polynomial terms up to the second order, which consists of $10$ basis: $\begin{bmatrix}1,x(t),y(t),z(t),x^{2}(t),x(t)y(t),x(t)z(t),y^{2}(t),y(t)z(t),z^{2}(t)\end{bmatrix}$. The training data is injected with 5% noise.
[…]
Figure 3 shows the prediction and the UQ result with different models. The dashed vertical line at $t=3.0s$ in the plot marks the switch between in-sample prediction and out-of-sample prediction, i.e., training data are the first 3,000 time steps, and we further perform 2,500 time steps of out-of-sample prediction. Figure 3 (a) illustrates the results in which the training data is injected with 5% noise. The predictions of the trajectories are based on the results in Table 2. It is observed that for in-sample prediction, LAPD and SINDy performs much better than UQ-SINDy and that LAPD prediction aligns better with the true trajectory compared to SINDy considering the out-of-sample prediction. The 95% confidence interval does not cover the true trajectory well for out-of-sample prediction since the prediction deviation accumulates with long predictions being made by solving the initial value problem. As shown in 3 (b), when the data is noise-free, the UQ of LAPD performs well. This result indicates that in the face of rather high noise, in order to provide good prediction and UQ, the time length of the prediction shouldn’t be so long to mitigate accumulated error. One may consider collecting new data in the later time step and setting it as a new initial condition instead of still using the old initial value.
Figure 3: (a) Trajectory prediction and confidence interval with SINDY, UQ-SINDy and LAPD on Lorenz system. Training data is injected with 5% noise. The dashed vertical line marks the transition from in-sample prediction to out-of-sample prediction. The confidence interval (95%) for LAPD is illustrated as green shaded area. (b) Trajectory prediction and confidence interval with LAPD on the Lorenz system. Training data is noise-free.
Table 2: Learned coefficients with different models on Lorenz System with 5% noise.
Basis True SINDy UQ-SINDy LAPD
X Y Z X Y Z X Y Z X Y Z
1 - - - - 5.672e-1 -5.545e-1 1.855e-2 4.449e-2 -2.583e-1 - - -
x -1.000e1 2.800e1 - -1.016e1 2.714e1 - -1.096e-1 2.250e1 -3.965e-3 -9.942e0 2.797e1 -
y 1.000e1 -1.000e0 - 1.014e1 -6.566e-1 - 7.420e-1 5.070e-1 -3.649e-3 9.981e0 -9.111e-1 -
z - - -2.667e0 - - -2.655e0 2.712e-2 8.359e-3 -1.377e0 - - -2.681e0
$x^{2}$1.728e-3 -6.264e-3 1.707e-1 - - -
$xy$ 1.000e0 - - 1.001e0 -2.500e-2 -4.793e-4 8.803e-1 - - 1.001e0
$xz$1.000e0 - - -9.856e-1 - -3.124e-1 -8.601e-1 2.113e-3 - -1.007e0 -
$y^{2}$ 1.337e-2 3.945e-3 -1.076e-2 - - -
$yz$ 3.202e-1 -1.363e-2 -5.477e-4 - - -
$z^{2}$ 1.384e-3 3.450e-4 -4.770e-2 - - -
Error Bar - - 5.267e4 1.435e-3
MSE - 2.648e0 7.994e1 1.024e0
AIC - 2.940e3 1.320e4 8.386e1
4.3 Burgers’ Equation
Burgers’ equation is a non-linear PDE system that has been widely investigated due to its versatility in modeling various physical processes. It serves as a simplified framework for understanding phenomena such as fluid dynamics, gas dynamics, and traffic flow. Its mathematical structure, which balances nonlinear advection and viscous diffusion, makes it a widely used test case for developing and validating numerical methods. The equation is given by:
$\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-\nu\frac{\partial^{2}u}{\partial x^{2}}=0,$ (15)
where $u(x,t)$ represents the velocity field, and $\nu$ denotes the viscosity coefficient. In this experiment, $\nu$ takes the value of 0.1. The training data for Burgers’ Equation is generated with 256 different positions spaced equally with $x\in[-8,8]$, and 101 different time steps spaced equally with $t\in[0,10]$. The candidate library consists of 10 terms with polynomial order up to 2 with the basis being $[1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. Initial condition is set to $u(0,x)=exp(-(x-3)^{2}/2)$. The training data is injected with 0.2% noise.
[…]
Table 3: Learned coefficients with different models on Burgers’ Equation with 0.2% Noise
Basis Burgers’ Equation
True SINDy UQ-SINDy LAPD(SGLD) LAPD(cyclical SGMCMC) LAPD(reSGLD)
$1$ 1.315e-3 - - -
$u$ 8.660e-2 5.908e-3 - - -
$u_{x}$2.195e-1 -1.363e-2 - - -
$u_{xx}$ 1.000e-1 5.431e-2 5.431e-2 9.161e-2 9.276e-2 9.274e-2
$u^{2}$1.960e-1 -3.109e-2 - - -
$uu_{x}$1.000e0 -2.476e-1 -9.494e-1 -9.642e-1 -9.908e-1 -9.915e-1
$uu_{xx}$ 1.921e-1 3.481e-2 - - -
$u_{x}^{2}$ 1.596e-2 - - -
$u_{x}u_{xx}$4.601e-2 - - -
$u_{xx}^{2}$ 7.756e-3 - - -
Error Bar - - 3.503e-2 2.983e-2 1.402e-4 1.218e-4
MSE - 1.448e-3 4.066e-4 9.386e-5 1.473e-5 1.421e-5
AIC - -1.690e5 -2.019e5 -2.398e5 -2.877e5 -2.886e5
Figure 4: Solution $u$ Contours of the identified Burgers’ Equation compared with ground truth (True $u$). Top: LAPD with default reSGLD as sampler Bottom: SINDy. Left Column: Contours of True $u$, Middle Column: Contours of predicted $\hat{u}$, Right Column: Contours of Error.
Figure 5: (a)Posterior distributions of identified model coefficients for Burgers’ Equation with different Langevin MCMC samplers adopted in LAPD. (b) An illustration of Error Bar (plotted in log scale) guiding threshold selection, $(0.5-0.7)$, with convection-diffusion equation.
4.4 Convection-Diffusion Equation
Convection-diffusion equation is a widely studied partial differential equation that describes the transport of a scalar quantity under the combined effects of convection and diffusion. It plays a critical role in modeling processes such as heat transfer, pollutant dispersion in fluids, and mass transfer in chemical engineering. The balance between convective transport, which moves the quantity with a flow, and diffusive transport, which acts to spread the quantity, makes it a fundamental equation for analyzing transport phenomena. The equation is expressed as:
$\frac{\partial u}{\partial t}+c\frac{\partial u}{\partial x}-D\frac{\partial^{2}u}{\partial x^{2}}=0,$ (16)
where $u(x,t)$ represents the transported quantity, $c$ is the convection velocity, and $D$ is the diffusion coefficient. The training data for Convection-Diffusion Equation is generated with 201 different positions spaced equally with $x\in[0,20]$, and 501 different time steps spaced equally with $t\in[0,5]$. The candidate library consists of 10 terms with polynomial order up to 2 with the basis being $[1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. The training data is injected with 0.1% noise.
In this experiment, we first discover the configuration of the thresholding settings. Figure 5(b) illustrates how the Error Bar value changes with different threshold settings for the Convection-Diffusion equation. The clear U-shaped curve indicates an optimal balance between model complexity and accuracy. At lower thresholds, the model includes too many terms, while at higher thresholds, essential terms are excluded, driving the model to learn poorly with higher standard deviation for coefficients. The Error Bar reaches its minimum in the range of 0.5 to 0.9, providing a data-driven approach to threshold selection in the sequential thresholding process. Subsequently, Table 4 compares the performance of SINDy, UQ-SINDy, and the proposed LAPD method in identifying the convection-diffusion equation under 0.1% noise. The LAPD framework achieves the closest alignment with the true coefficients (e.g., $u_{x}:-1.000$, $u_{xx}:0.960$), while other methods exhibit significant deviations or spurious terms (e.g., erroneous $uu_{x}$ or $u_{xx}^{2}$ coefficients). LAPD also demonstrates superior robustness, with the lowest MSE ($2.212\times 10^{-3}$), minimal posterior uncertainty ($8.613\times 10^{-4}$), and optimal AIC ($-6.157\times 10^{5}$), underscoring its efficacy in sparse, noise-resilient dynamics discovery.
Figure 6: Solution $u$ Contours of the identified convection-diffusion Equation using LAPD compared with ground truth (True $u$). Top: Initial condition IC1, Bottom: Initial condition IC2. Left Column: Contours of True $u$, Middle Column: Contours of predicted $\hat{u}$, Right Column: Contours of Error.
Table 4: Learned coefficients with different models on Convection-Diffusion Equation with 0.1% noise
Basis Convection-Diffusion Equation
True SINDy UQ-SINDy LAPD(SGLD) LAPD(cyclical SGMCMC) LAPD(reSGLD)
$1$ 4.090e-4 - - -
$u$ 4.051e-2 - - -
$u_{x}$1.000e0 -1.265e0 -1.200e0 -9.985e-1 -1.012e0 -1.000e0
$u_{xx}$ 1.000e0 - 4.686e-1 9.514e-1 9.551e-1 9.600e-1
$u^{2}$6.343e-1 - - -
$uu_{x}$ 1.291e0 1.283e-0 - - -
$uu_{xx}$ 5.544e-1 1.460e0 - - -
$u_{x}^{2}$ 1.148e0 - - -
$u_{x}u_{xx}$8.611e-3 - - -
$u_{xx}^{2}$ 3.627e-1 9.536e-1 - - -
Error Bar - - 1.123e1 5.141e-3 1.486e-3 8.613e-4
MSE - 5.886e-2 5.782e-2 2.327e-3 2.278e-3 2.212e-3
AIC - -2.852e5 -2.870e5 -6.106e5 -6.127e5 -6.157e5
5 Experiments with Active Learning
Given the limited complexity of the sample size inherent to AL settings, we employ the Metropolis adjusted Langevin algorithm (MALA) with exact gradients via LAPD, which serves as a computationally efficient single-chain alternative to reSGLD.
5.1 Lotka-Volterra System with Active Learning
We evaluate the LAPD framework on the Lotka-Volterra system, following the procedure outlined in Algorithm 1. The data pool for this experiment, denoted as $\mathcal{D}$, consists of $N=10,000$ points sampled from the states generated in the 50,000 time steps. The data simulating procedure is the same as before, except for $50,000$ time steps of states are collected, with $\Delta t=5\cdot 10^{-3}$. Furthermore, 5% noise is also injected during experiments. With $\mathcal{D}$ being randomly sampled from the simulated data, it can be viewed as an irregularly sampled data set. In this synthetic experiment setting, the acquisition of the time derivative is carried out using the original regularly sampled data. However, in realistic conditions, the acquisition of the response variable (the time derivative) is computationally expensive. Here $\lambda=0,\alpha=0.5$ are adopted.
Figure 7: Illustration of active learning for the LAPD method on the Lotka–Volterra system. (a) All potential sample points (light gray) and the initial 20 randomly chosen training points (dark gray) for the first round. Red points show points added over the four following iterations using the Uncertainty + Space-Filling acquisition function, whereas blue points reflect the purely Uncertainty-based selection. (b) Log-scale plot of the error bar for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added.
Training begins with an initial set of 20 randomly selected points. In each subsequent round, 10 additional points are selected via the specified acquisition functions and added to $\mathcal{M}$. Based on the acquisition function of Uncertainty+Space Filling, the threshold Tol is set to 0.8, and experiments with other acquisition functions are carried out for the same number of rounds. As depicted in Figure 7(a), combining space filling and uncertainty criteria yields a more uniform spatial distribution of the sampled points, while relying solely on uncertainty concentrates the sampling in regions of high state magnitude. Clustered data in one area have a higher chance of model overfitting and lack representativeness of the whole dataset. As shown in Figure 7(b) and Table 5, although the error bar in the round with 60 points does not differ much ($\Delta=6\cdot 10^{-3}$ on linear scale), the learned system combining Uncertainty and Space Filling yields a learned system more closely aligned with the true dynamics. Furthermore, this combined strategy illustrates the efficiency of systematic AL over random sampling. For the error bar to drop below $10^{-2}$, the random sampling strategy requires 70 data, while the combined strategy uses 30 data, reducing required measurements by around 60%. AL acquisition functions enhance the informativeness and representativeness of selected points, other than training efficiency, under the same round of training; it also improves alignment with the ground truth dynamics. As shown in Table 5, compared to randomly adding points, the Uncertainty and Uncertainty+Space-Filling strategies more accurately preserve the nonlinear structure of the ground truth, capturing both growth and predator–prey interactions without extra terms.
Table 5: Comparison of true system (Lotka Volterra) and identified systems for different strategies with 60 training points
System $x_{t}$ $y_{t}$
True system $x_{t}=x-0.1xy$ $y_{t}=-1.5y+0.075xy$
Random $x_{t}=0.181+0.992x-0.099xy$ $y_{t}=-0.311-1.449y+0.073xy$
Uncertainty $x_{t}=0.902x-0.091xy$ $y_{t}=-1.339y+0.067xy$
Uncertainty+Space-Filling $x_{t}=0.998x-0.096xy$ $y_{t}=-1.423y+0.070xy$
5.2 Burgers’ Equation with Active Learning
We next evaluate LAPD on Burgers’ Equation. The original data here distributed around 4001 positions equally spaced in [-8,8] and 1001 different time steps equally spaced in [0,10]. The initial condition is set to the same as in Section 4.3. Among the original data, 3 different time steps are selected as the training time step. Here, we chose 1s,5s, and 8s, and actually, this time step can be chosen randomly, and even the whole training dataset can be chosen randomly throughout the whole dataset. Under this scenario, the calculation of partial derivatives with respect to $t$ can be difficult to obtain, hence in need of active learning. Again, in this synthetic data scenario, we calculate $\mathbf{U}_{t}$ based on the original data. The candidate library adopted in this experiment is the same as in Section 4.3 with 10 terms $[1,u,u_{x},u_{xx},u^{2},uu_{x},uu_{xx},u_{x}^{2},u_{x}u_{xx},u_{xx}^{2}]$. Here we inject i.i.d. noise on $\dot{\mathbf{U}}$ following $lognormal(0,0.1)$.
Training begins with an initial set of 20 randomly selected data. In each subsequent round, 10 additional data are selected via the specified acquisition functions and added to $\mathcal{M}$. After the first round of training with the initial data, the minimum $d(\mathbf{u})$ is $6.769\times 10^{-11}$, while the maximum $d(\mathbf{u})$ is $1.443$, indicating a significant relative difference of more than $10^{10}$, so density adjustment is performed. Here, we adopt $\lambda=0.5,\alpha=0.3$. As shown in Figure 8, AL with a hybrid acquisition function performs better, requiring fewer training points. The threshold Tol is set to 0.3, and experiments with other acquisition functions are carried out for the same number of rounds as the uncertainty+space-filling acquisition function. The hybrid method identifies the correct candidate functions with 50 points, while the Uncertainty method and the Random method are left with only 4 and 7 basis functions, respectively. For the error bar to drop below $10^{-2}$, the random sampling strategy requires 90 data, whereas the hybrid strategy only requires 50 data, reducing the required measurements by 44%. As shown in Table 6, the random sampling approach results in an equation contaminated with extraneous terms such as $u,u^{2},uu_{x},u_{x}u_{xx}$, which deviate significantly from the true system. This suggests that randomly selected training points fail to capture the underlying physics, leading to overfitting or the inclusion of spurious terms. On the other hand, uncertainty-based sampling shows improvement by reducing the number of incorrect terms. However, the coefficient estimates remain suboptimal, particularly for the nonlinear interaction term $u\frac{\partial u}{\partial x}$. This indicates that an uncertainty-only acquisition function may still suffer from biased selection, potentially over-exploring regions with high variance while neglecting global representativeness. The hybrid uncertainty+space-filling strategy demonstrates the most accurate identification of the governing equation. The identified terms match the true system structure, and the coefficient values closely approximate their theoretical counterparts. This highlights the advantage of balancing exploration (uncertainty) and diversity (space-filling) when selecting training points, leading to more robust equation discovery with fewer data points.
Figure 8: Active Learning on Burgers’ Equation: Log-scale plot of the error bar for Random, Uncertainty, and Uncertainty + Space-Filling acquisition functions as training points are incrementally added.
Table 6: Comparison of true system (Burgers’ Equation) and identified systems for different strategies with 70 training data
System $\frac{\partial u}{\partial t}$
True system $\frac{\partial u}{\partial t}+u\frac{\partial u}{\partial x}-0.1\frac{\partial^{2}u}{\partial x^{2}}=0$
Random $\frac{\partial u}{\partial t}-0.054u-0.075\frac{\partial u}{\partial x}+0.147\frac{\partial^{2}u}{\partial x^{2}}+0.944u^{2}+0.060u\frac{\partial u}{\partial x}+0.115\frac{\partial u}{\partial x}\frac{\partial^{2}u}{\partial x^{2}}=0$
Uncertainty $\frac{\partial u}{\partial t}+0.14\frac{\partial u}{\partial x}+0.714u\frac{\partial u}{\partial x}-0.073\frac{\partial^{2}u}{\partial x^{2}}=0$
Uncertainty+Space-Filling $\frac{\partial u}{\partial t}+0.971u\frac{\partial u}{\partial x}-0.090\frac{\partial^{2}u}{\partial x^{2}}=0$
6 Conclusion and Discussion
This work presents a novel Lagevin-assisted Bayesian active learning framework, LAPD, which integrates reSGLD with AL to discover governing equations from experimental data. Our comprehensive evaluation across diverse dynamical systems demonstrates that incorporating reSGLD sampling within the physical discovery process provides significant advantages over traditional approaches, especially in the face of noise. By simultaneously operating multiple Markov chains at different temperatures, LAPD achieves a faster exploration of the parameter space, effectively navigating the complex landscapes of nonlinear dynamical systems. Traditional sparse regression methods for physical discovery (like SINDy) provide point estimates of system coefficients but offer limited insight into model confidence intervals. In contrast, LAPD offers robust UQ for both model parameters and system trajectories. Moreover, while sparsity-inducing priors (as in UQ-SINDy) can theoretically promote sparse solutions, our experiments improved robustness and accuracy when identifying system dynamics and estimating the uncertainties. Furthermore, our AL extension introduces a hybrid acquisition function that balances uncertainty reduction with domain coverage. This approach significantly outperforms both random sampling and pure uncertainty-based acquisition, reducing measurement requirements by 44-57% across test cases.
While the LAPD demonstrates robust performance across our test cases, several methodological challenges remain. First, the computational cost of reSGLD sampling exceeds that of deterministic optimization approaches. For very large systems or real-time applications, this overhead may become prohibitive. Algorithmic optimizations like adaptive temperature scheduling or parallel chain implementation could address this limitation. Second, our current implementation determines the initial library of candidate functions a priori. For complex systems where appropriate basis functions are unknown, this presents a practical challenge. Future work could integrate automatic basis function discovery, perhaps by incorporating neural networks to learn transformed coordinates where dynamics admit sparse representations.
In summary, LAPD represents a significant step toward robust, uncertainty-aware physical discovery from limited data. By bridging Bayesian inference, stochastic sampling, and AL, it addresses fundamental challenges in data-driven science. As experimental data in many fields remain costly and imperfect, methods like LAPD that maximize information extraction while quantifying uncertainty will become increasingly valuable for scientific discovery.